Suppose that you flip a coin 12 times. What is the probability that you achieve at least 4 tails?
Answer by greenestamps(13203)



# of ways of getting 0 tails: C(12,0) = 1
# of ways of getting 1 tails: C(12,1) = 12
# of ways of getting 2 tails: C(12,2) = 66
# of ways of getting 3 tails: C(12,3) = 220

# of ways of getting less than 4 tails: 1+2+66+220 = 299

# of ways of getting at least 4 tails: 2^12 - 299 = 4096-299 = 3797

Probability of getting at least 4 tails: 3797/4096 = 92.7%